# miniCodeProps: a Minimal Benchmark for Proving Code Properties

### v1

@@ Abstract
Neural networks have shown initial promise in automating mathematical theorem proving in proof assistants such as Lean. The same proof assistants can be used to verify the correctness of code by pairing code with specifications and proofs that the specifications hold. Automating the writing of code, specifications, and proofs could lower the cost of verification, or, ambitiously, enable a machine learning system to output provably correct code. However, it remains unclear whether current neural theorem provers can automatically verify even relatively simple programs. We present miniCodeProps, a benchmark of 177 program specifications in the Lean proof assistant, aimed at the subproblem of automatically generating a proof for a provided program and specification. miniCodeProps contains specifications about simple, self-contained programs (e.g., lists, natural numbers, binary trees) with varied proof difficulty. Despite its simplicity, miniCodeProps is challenging for current LLM-based provers, which succeed in proving about 25 percent of the specifications. We publicly release miniCodeProps as a benchmark for furthering automated theorem proving in the context of formally verified code.
@@ 1 Introduction
Writing code that meets a specification is a desirable, yet difficult task. In safety-critical contexts, users require stronger safeguards than code reviews and test cases provide. To address this need, the formal methods community has produced a variety of tools for specifying and proving properties of code [1, 2, 3, 4, 5]. One approach uses interactive theorem provers (ITPs) such as Isabelle [6], Coq [7], and Lean [8]. In an ITP, a user writes code, a specification, and a proof that the specification holds (Figure 1). The ITP checks each step of the proof as it is written, and a complete proof means that the specification is guaranteed to hold. A complete proof thus gives guarantees on the underlying code, which is valuable in practice. For instance, the Lean proof assistant was recently used to verify properties of Amazon’s Cedar Policy language. [^1] https://github.com/cedar-policy/cedar-spec More speculatively, future code generation models could in principle output provably correct code if they are able to write code, specifications, and proofs [9].
@@
Despite its promise, interactive theorem proving remains a difficult task, and better automation–be it writing specifications, proofs, or combinations thereof–could help make it easier to verify code. Recently, machine learning techniques based on large language models (LLMs) have shown promise in automatically generating code [10], proving mathematical theorems in ITPs [11], and generating proofs of properties from verification projects [12]. However, evaluating the capabilities of LLMs for program verification remains a challenge. For example, benchmarks often evaluate across projects with complex dependencies such as the CompCert compiler or Archive of Formal Proofs [13, 12], making it difficult to isolate and reason about a model’s core capabilities and weaknesses.
@@
Figure 1: Formal verification of code in an interactive proof assistant (here, Lean) consists of writing (i) code (e.g., an implementation of a Tree), (ii) a property that you want to verify (e.g. that the size of a balanced tree is odd), and (iii) a proof that the code satisfies the property. The underlying language verifies that the proof is correct, providing certainty that the property holds. miniCodeProps focuses on the important subproblem of proof generation: given a property and associated code, a model must generate a proof (e.g., shown here in green) that the property holds. miniCodeProps contains 177 properties about simple and self-contained code blocks, and measures a range of proving abilities.
@@
To address this need, we introduce miniCodeProps, a benchmark for evaluating the ability to prove properties of relatively simple, self-contained programs in the Lean proof assistant. miniCodeProps contains 177 specifications about simple, self-contained programs (involving e.g., lists, natural numbers, binary trees) with varied proof difficulty, sourced by translating Haskell programs from Tons of Inductive Programs [14] into Lean 4. Despite its simplicity, miniCodeProps is challenging for current LLM-based provers. For example, our best baseline approach based on GPT-4 proved very few specifications requiring proofs longer than a few lines. We publicly release miniCodeProps as a benchmark for furthering automated theorem proving in the context of formally verified code. [^2] Our benchmark and baselines can be found here: https://huggingface.co/datasets/elohn/miniCodeProps
@@ 2 Related Work
There is a rich history of developing automation for formal verification of code; see [4] for a survey. For machine learning in interactive theorem proving, many methods that preceded large language models focused on Coq, such as ProverBot9000 [26], ASTactic [13], TacTok [27], Diva [28], and Passport [29]. Recent work with LLMs includes exploring prompting strategies in Coq [30], and Baldur [12] which explores proof generation and repair in Isabelle’s Archive of Formal Proofs. We are not aware of machine learning-based proof automation targeting program verification in Lean, despite the activity in automated mathematical proving in Lean. One of our motivations is bridging this gap. Finally, some methods such as COPRA [24] aim to be language-agnostic; miniCodeProps could participate in evaluating and developing such systems.
@@ 3 Benchmark Contents
We create miniCodeProps by translating programs from Tons of Inductive Programs (TIP) [14] (https://github.com/tip-org/benchmarks) from Haskell into Lean 4. We manually translated code from three files in TIP that contain a mix of function definitions, propositions describing the results of calling those functions, and termination lemmas used in the function definitions. In total, this yields 177 Lean 4 theorem statements.
@@ 3.1 Collection Process
The properties in TIP describe induction in both programming and purely mathematical contexts. We first filtered out files that contained primarily mathematical induction, then translated the remaining Haskell code into Lean 4. We then associated each property in the translated files with dependencies and metadata useful for different evaluation modes of the benchmark.
@@ 3.2 Source Code Composition
3.2.1 Functions
Most of the translated functions operate on linked lists, while the rest involve natural numbers and binary trees. With a few notable exceptions, the functions perform conceptually simple operations such as filtering, returning the last element, and counting the elements of a list. The more complicated functions are increasingly esoteric sorting functions. In total, miniCodeProps is derived from 76 TIP functions. A breakdown of the primary subject of each function is provided in Table 1.
Numbers Lists Trees Heaps
Medley Functions 0 15 2 0
Medley Properties 21 64 1 0
@@ 3.2.1 Functions
Termination Properties 3 22 0 3
Sorting Properties 0 63 0 0
Table 1: miniCodeProps Composition We classified each function and property in the Lean source files by their primary subject matter. We also separate these numbers by the splits of miniCodeProps: Medley, Termination lemmas, and Sorting algorithms. Termination and Sorting algorithm functions are grouped because the properties in both splits describe the same set of functions.
3.2.2 Properties
The properties to be proven in miniCodeProps express intuitively correct properties of the function(s) being described. See Section 3.4 for examples, and Table 1 for a breakdown by subject type.
3.2.3 Termination Lemmas
In Lean 4, recursive functions must be paired with a proof of termination. While Lean automatically infers the proof in simple cases, more complicated recursive calls require the user to explicitly prove termination. As the vast majority of our function definitions are recursive, our benchmark includes 28 lemmas that support 4 termination proofs of nonstandard sorting algorithms from TIP. These lemmas themselves represent practically useful and highly nontrivial properties of code.
3.3 Dataset Splits
To improve understanding of the precise capabilities of theorem proving agents when evaluated on miniCodeProps, we section the 177 properties into 3 data splits: Medley, Termination lemmas, and Sorting algorithms. Medley properties are named “prop_n” for $n\in[1,86]$, and tend to involve simple functions on numbers and lists. Sorting algorithm properties are those remaining theorems named beginning with “prop”, and the remainder are Termination lemma properties.
The benchmark is formatted as a jsonlines file with the information in Table 2 per line. The “full_name” property is used to determine the split the property belongs to.
@@ 3.4 Qualitative Examples
We present several examples from the benchmark and describe what makes them interesting and difficult.
3.4.1 Zip over Concatenation
@@
This property is intuitively correct, but requires some intuition about proofs of recursive structures to correctly prove in Lean. The following is an incorrect proof generated by GPT4. It contains one semantic error (failing to generalize $xs$ in the inductive hypothesis) and one misuse of Lean internals; the ‘cases’ keyword should be used instead of ‘match.’
induction ys with
| nil =>
simp [zip’, List.take, List.drop]
| cons y ys ih => match xs with
| [] => simp [zip’, List.take,
List.drop, List.length]
| cons x xs =>
unfold zip
unfold List.take at ih
unfold List.drop at ih
simp [List.length]
rw [ih]
The following is a correct proof of prop_84 generated by GPT4. There are several unnecessary definitions passed to the calls to the simplifier (the ‘simp’ tactic). However, in both the correct and incorrect proofs, GPT4 demonstrates the main ideas of the proof: induct on $ys$, pair off the first elements of $xs$ and $ys$, then apply the inductive hypothesis to what remains.
@@
3.4.2 Non-standard Insertion Increments Length
@@
The following is an incorrect proof of ‘prop_15’ by GPT4. Interestingly, GPT4’s main failing in this case is syntactic; the proof can be manually corrected by using the correct ‘by_cases‘ tactic in place of ‘byCases’ unrecognized by Lean 4 and manual removal of the following case matching (the ‘| in…’ statements), at which point the proof is completed at the ‘rw [ih]’ step.
induction xs with
| nil =>
simp [ins, List.length, Nat.succ]
| cons y ys ih =>
simp [ins]
byCases h : x < y
| inl h_lt =>
simp [h_lt, List.length, Nat.succ]
| inr h_geq =>
simp [h_geq]
rw [ih]
simp [Nat.succ_assoc]
As one might expect from the “nearly correct” proof above, GPT4 also produces a correct proof. This proof is remarkable for its simplicity. Unlike the previous correct proof there is little wasted space, a feature often sought after by human proof experts. GPT4 demonstrates understanding of how to work with the ‘ins’ function via the ‘split_ifs’ tactic, and the usage of the inductive hypothesis only in the second branch where it is necessary.
@@ 3.4.3 Sorting Properties
The properties above express the following intuitive properties: heapsort orders list elements, preserves the list size, returns a permutation of the original list, and returns the same result as insertion sort. These properties are repeated for each sorting algorithm. However, given the complexity of the sorting functions involved, we expect that proving such properties will be very difficult. miniCodeProps also includes termination lemmas such as ‘numElem_merge_branches_lt’ with the human-produced proof removed. An example of an incorrect proof of one of the sorting properties produced by GPT4 can be found in Appendix A.
3.5 Availability and Usage
We have published our benchmark as a public dataset on Huggingface, including a link to our example benchmarking code. The TIP repo our benchmark is based on uses the BSD-3-Clause license, permitting re-use and modification with license redistribution; we therefore use the same. Given the information in Table 2, there are many possible tasks relevant for testing an agent’s proof generation capabilities. We propose the following as interesting evaluation modes.
  1. 1.
    Proof State only: Past work has trained next-step tactic prediction models taking only the proof state provided by the proof assistant as input [16, 19, 21, 35]. This input is justifiable in contexts such as mathematical theorem proving where most proofs involve a shared mathematical library, because agents will be exposed to most of the definitions and lemmas used in the goal during training. We expect this input to be insufficient for many proofs about programs, as the programs referenced in the proof state are only described in the file context. However, this mode will test an agent’s ability to apply reasonable proof techniques when lacking knowledge of certain terms in the proof state.
  2. 2.
    Proof State + Dependencies: In this mode, agents have access to all newly written (i.e. not from Lean core or Mathlib) code and lemmas required to define the given code property. This can be accomplished with the “deps” and “proof_state” fields of each benchmark entry.
  3. 3.
    Proof State + Full File Context: In this mode, the full contents of the file(s) prior to the property definition are made available. Similar to Proof State + Dependencies mode, this input form gives an agent access to all definitions in the proof state. However, full file context mode requires the agent to filter its input for useful context, and additionally may give the agent access to potentially useful lemmas defined in the file that were not necessary to define the property. Further, this mode most closely matches what a human expert must do when proving such properties.
  4. 4.
    Feedback Refinement: In this mode, error feedback from (potentially partial) proof suggestions is used as input to continued attempts at proving the property until a given number of attempts is reached. In practice, Lean is intended to be used interactively. Even for simple proofs, it may take human experts several iterations of error feedback from Lean to construct a correct proof. Enabling this feedback loop for theorem proving agents allows more fair comparison of current neural theorem proving approaches with human capabilities.
@@ 4 Baselines
We use two main approaches: full proof generation via few-shot prompting a language model and next-step tactic generation. In the full-proof context, the model generates one or more potential proofs that are verified by the proof checker (in our case, the Lean 4 kernel). For next-step tactic prediction, we follow the common practice of taking the proof state (and optionally file context) as input and returning suggestions for the next tactic to use in the proof. Each suggestion is then given as input to the Lean 4 kernel, and the resulting proof state is used to prompt the model for the next tactic.
Experimental Setup: Next-Step Tactic Generation
Our experiments with next-step tactic generation used the LLMStep [35] framework, modified to return cumulative log probabilities in addition to next step suggestions from each backend language model. The Pythia2.8B experiments were run with proof state only as input to match the model’s fine-tuning set. The remaining models received proof state + dependencies as input. These experiments were run on a single A100 GPU running Ubuntu 20.04 with 30 CPU cores rented from LambdaLabs.
Experimental Setup: Full Proof Generation
We employed a few-shot prompting approach. Our prompt contained examples of code, properties, and proofs of properties not seen in the benchmark. As such, the evaluation mode used was proof state + dependencies, although in practice we opted to use only the dependencies and the theorem statement itself. All experiments were performed using the OpenAI chat completions API with GPT4-turbo using pass@8.
Model Medley (Easy) Termination (Med.) Sorting (Hard)
llmstep+Pythia2.8B [35] 44/86 1/28 0/63
llmstep+Llemma-7B [25] 46/86 2/28 0/63
llmstep+ntp-context-1.3B [36] 38/86 0/28 0/63
GPT-4-turbo 44/86 1/28 0/63
Table 3: miniCodeProps results. Number of specifications proven when applying next-step tactic generation with LLMStep and full proof generation with GPT4 to the problem of verifying program specifications. The Medley section contains mostly of specifications that can be proven in several lines. Proofs of the sorting algorithm properties and termination lemmas are expected to require at least tens of lines and hours of programmer effort.

### v2

@@ Abstract
AI agents have shown initial promise in automating mathematical theorem proving in proof assistants such as Lean. The same proof assistants can be used to verify the correctness of code by pairing code with specifications and proofs that the specifications hold. Automating the writing of code, specifications, and proofs could lower the cost of verification, or, ambitiously, enable an AI agent to output safe, provably correct code. However, it remains unclear whether current neural theorem provers can automatically verify even relatively simple programs. We present miniCodeProps, a benchmark of 201 program specifications in the Lean proof assistant, aimed at the subproblem of automatically generating a proof for a provided program and specification. miniCodeProps contains specifications about simple, self-contained programs (e.g., lists, natural numbers, binary trees) with varied proof difficulty. Despite its simplicity, miniCodeProps is sufficient to break current LLM-based provers, with state-of-the-art methods showing promise on the easy properties in miniCodeProps, yet failing to prove nearly all of the medium and hard properties. We publicly release miniCodeProps as a benchmark for furthering automated theorem proving in the context of formally verified code. [^1] Our benchmark and evaluation code can be found here: https://github.com/cmu-l3/minicodeprops-eval
@@ 1 Introduction
Writing code that meets a specification is a desirable, yet difficult task. In safety-critical contexts, users require stronger safeguards than code reviews and test cases provide. To address this need, the formal methods community has produced a variety of tools for specifying and proving properties of code [1, 2, 3, 4, 5]. One approach uses interactive theorem provers (ITPs) such as Isabelle [6], Coq [7], and Lean [8]. In an ITP, a user writes code, a specification, and a proof that the specification holds (Figure 1). The ITP checks each step of the proof as it is written, and a complete proof means that the specification is guaranteed to hold. A complete proof thus gives guarantees on the underlying code, which is valuable in practice. For instance, the Lean proof assistant was recently used to verify properties of Amazon’s Cedar Policy language [9]. Ambitiously, future code generation models could also output safe, provably correct code if they are able to write code, specifications, and proofs [10].
@@
Despite its promise, interactive theorem proving remains a difficult task, and better automation–be it writing specifications, proofs, or combinations thereof–could help make it easier to verify code. Recently, machine learning techniques based on large language models (LLMs) have shown promise in automatically generating code [11], proving mathematical theorems in ITPs [12], and generating proofs of properties from verification projects [13]. However, evaluating the capabilities of LLMs for program verification remains a challenge. For example, benchmarks often evaluate across projects with complex dependencies such as the CompCert compiler or Archive of Formal Proofs [14, 13], making it difficult to isolate and reason about a model’s core capabilities and weaknesses. Rather than evaluating on complex projects, our goal is to find a minimal set of meaningful program properties that breaks current LLM-based provers, and thus requires new methods or models to solve.
@@
Figure 1: Formal verification of code in an interactive proof assistant (here, Lean) consists of writing (i) code (e.g., an implementation of a Tree), (ii) a property that you want to verify (e.g. that the size of a balanced tree is odd), and (iii) a proof that the code satisfies the property. The underlying language verifies that the proof is correct, providing certainty that the property holds. miniCodeProps focuses on the important subproblem of proof generation: given a property and associated code, a model must generate a proof (e.g., shown here in green) that the property holds. miniCodeProps contains 201 properties about lightweight, self-contained code blocks, and measures a range of proving abilities.
@@
To this end, we introduce miniCodeProps, a benchmark for evaluating the ability to prove properties of relatively simple, self-contained programs in the Lean proof assistant. miniCodeProps contains 201 specifications about simple, self-contained programs (involving e.g., lists, natural numbers, binary trees) with varied proof difficulty, sourced by translating Haskell programs from Tons of Inductive Programs [15] into Lean 4. Despite its simplicity, miniCodeProps is challenging for current LLM-based provers. For example, our best baseline approach based on GPT-4 proved very few specifications requiring proofs longer than a few lines. We publicly release miniCodeProps as a benchmark for furthering automated theorem proving in the context of formally verified code.
@@ 2 Related Work
There is a rich history of developing automation for formal verification of code; see [4] for a survey. For machine learning in interactive theorem proving, many methods that preceded large language models focused on Coq, such as ProverBot [27, 28], ASTactic [14], TacTok [29], Diva [30], and Passport [31]. Recent work with LLMs includes exploring prompting strategies in Coq [32], and Baldur [13] which explores proof generation and repair in Isabelle’s Archive of Formal Proofs. We are not aware of machine learning-based proof automation targeting program verification in Lean, despite the activity in automated mathematical proving in Lean. One of our motivations is bridging this gap. Finally, some methods such as COPRA [25] aim to be language-agnostic; miniCodeProps could participate in evaluating and developing such systems.
@@ 3 Benchmark Contents
We create miniCodeProps by translating programs from Tons of Inductive Programs (TIP) [15] (https://github.com/tip-org/benchmarks) from Haskell into Lean 4. We manually translated code from three files in TIP that contain a mix of function definitions, propositions describing the results of calling those functions, and termination lemmas used in the function definitions. In total, this yields 201 Lean 4 theorem statements. We specifically make miniCodeProps fairly small, since evaluating theorem proving models requires many calls to the language model (e.g., in a tree search or through repeated sampling), and is hence time-consuming and prohibitively expensive for large datasets.
@@ 3.1 Collection Process
The properties in TIP describe induction in both programming and purely mathematical contexts. Since our focus is programming, we filtered out files that contained primarily mathematical induction, then translated the remaining Haskell code into Lean 4. We associated each property in the translated files with dependencies and metadata useful for different evaluation modes of the benchmark.
@@ 3.2 Source Code Composition
miniCodeProps contains properties that we view as a “minimum level of competency” for automated neural theorem provers. We draw an analogy to an expert software engineer that spends time writing code in a complex repository, yet is expected to be capable of solving self-contained interview problems. miniCodeProps embodies the latter, which is absent from existing theorem proving benchmarks. Namely, rather than pulling code and properties from complex verification projects, miniCodeProps tests key capabilities with lightweight, self-contained code blocks and properties.
These fall into three categories: (1) intuitive properties of lists, trees, and natural numbers (medley), (2) lemmas that support termination proofs, which are an essential (and tedious) part of verification in an interactive theorem prover (termination); (3) properties of nonstandard sorting algorithms that require a deep understanding of the corresponding code (sorting). We provide more detail below.
Functions.
Most of the translated functions operate on linked lists, while the rest involve natural numbers and binary trees. With a few notable exceptions, the functions perform conceptually simple operations such as filtering, returning the last element, and counting the elements of a list. The more complicated functions are increasingly esoteric sorting functions. In total, miniCodeProps is derived from 95 TIP functions. A breakdown is provided in Table 1.
Numbers Lists Trees Heaps Difficulty
Medley Functions 0 26 10 0
Medley Properties 21 84 5 0 Easy
@@
Termination Properties 3 22 0 3 Medium
Sorting Properties 0 63 0 0 Hard
Table 1: miniCodeProps Composition We classified each function and property in the Lean source files by their primary subject matter. We also separate these numbers by the splits of miniCodeProps: Medley, Termination lemmas, and Sorting algorithms. Termination and Sorting algorithm functions are grouped because the properties in both splits describe the same set of functions. The properties are classified into easy (Medley), medium (Termination), and hard (Sorting) difficulty levels.
Properties.
The properties to be proven in miniCodeProps express intuitive properties of the function(s) being described. Properties in the medley classification typically require a standard induction argument along with the application of 1-2 library theorems. Properties of the nonstandard sorting algorithms are easy to state, yet are likely much more difficult prove due to the complexity of the sorting functions. See Section 3.3 for examples, and Table 1 for a breakdown by subject type.
Termination Lemmas.
In Lean 4, recursive functions must be paired with a proof of termination. While Lean automatically infers the proof in simple cases, more complicated recursive calls require the user to explicitly prove termination. As the vast majority of our function definitions are recursive, our benchmark includes 28 lemmas that support 4 termination proofs of nonstandard sorting algorithms from TIP. These lemmas represent highly nontrivial properties of code that, roughly speaking, may take hours of human effort to prove. Automating these proofs would be highly desirable in practice.
Dataset format.
The benchmark is formatted as a jsonlines file with the information in Table 2 per line. The “full_name” property is used to determine the split the property belongs to: medley properties are named “prop_n” for $n\in[1,86]$, sorting algorithm properties are those remaining theorems named beginning with “prop”, and the remainder are termination lemma properties.
@@ 3.3 Qualitative Examples
We study several benchmark examples and describe what makes them interesting and difficult.
Zip over concatenation.
@@
This property is intuitively correct, but requires some intuition about proofs of recursive structures to correctly prove in Lean. The following is a correct proof of prop_84 generated by GPT4. There are several unnecessary definitions passed to the calls to the simplifier (the ‘simp’ tactic). However, GPT4 demonstrates the main ideas of the proof: induct on $ys$, pair off the first elements of $xs$ and $ys$, then apply the inductive hypothesis to what remains.
@@
Many of the medley properties test similar abilities, which we view as bare minimum capabilities for a system aimed at automating theorem proving in the context of formally verified code.
Non-standard insertion increments length.
@@
The following is a correct proof generated by GPT-4. This proof is remarkable for its simplicity. Unlike the previous correct proof there is little wasted space, a feature often sought after by human proof experts. GPT4 demonstrates understanding of how to work with the ‘ins’ function via the ‘split_ifs’ tactic, and the usage of the inductive hypothesis only in the second branch where it is necessary.
@@
The properties above express the following intuitive properties: heapsort orders list elements, preserves the list size, returns a permutation of the original list, and returns the same result as insertion sort. These properties are repeated for each sorting algorithm. However, given the complexity of the sorting functions involved, we expect that proving such properties will be very difficult. An example of an incorrect proof of one of the sorting properties produced by GPT4 can be found in Appendix A.
Finally, miniCodeProps includes termination lemmas such as ‘numElem_merge_branches_lt’ with the human-produced proof removed. Some of these lemmas took hours of human effort to prove. Since proving termination is an essential part of human interactive theorem proving, it is highly likely that improved performance on these lemmas could translate into useful automation for practitioners.
Availability and usage.
We have published our benchmark as a public dataset on Huggingface, and released a repository for evaluating language models on miniCodeProps.
@@ 4 Baselines
We test two standard proving modes: full-proof generation and tactic-by-tactic generation. Since the underlying code block for a property is typically necessary to complete the proofs, we always provide the models with the code dependencies (stored as “deps” in our dataset, see Table 2).
For full-proof generation, the model generates one or more potential proofs that are checked by the proof checker (in our case, the Lean 4 kernel). For tactic-by-tactic generation, we follow the common practice of taking the proof state as input and returning suggestions for the next tactic to use in the proof. Each suggestion is then given as input to the Lean 4 kernel, and the resulting proof state is used to prompt the model for the next tactic. We test the de-facto standard algorithm for tactic generation, best-first search [17, 19, 22, 26]. Our publicly available evaluation code supports both modes.
Tactic-by-tactic generation. We use ntp-ctx-1.3B [37] since at the time of experimentation it is the only tactic generation model specifically trained to accept both the current proof state and an additional input context (in our case, the necessary underlying code for the property being proven). The ntp-ctx-1.3B model is a DeepSeek-Coder 1.3B language model fine-tuned on (context, proof state, next tactic) examples derived from Mathlib [37]. For the context, following a similar setup to [37] we provide a context containing the dependent code (“deps” in Table 2), the property being proven, and the history of generated tactics. We use the same input format and search settings as [37].
Full proof generation. We use GPT-4o, since at the time of experimentation it is a highly-performant general-purpose language model on mathematical and code tasks, and its model variants (e.g., previous models of GPT-4) are common baselines in neural theorem proving (e.g., [25, 37]). We provide a context containing the dependent code (“deps” in Table 2) and the property being proven. We measure pass@32, meaning that we generate 32 proof candidates (with default generation parameters), and consider a property to be successful when at least one of them passes the verifier.
Full proof generation with refinement. As a strictly stronger full proof generation baseline, we perform 2 rounds of an experimental “many-to-one” refinement technique using GPT-4o. Specifically, suppose that none of the 32 candidate proofs $y^{(1)},\ldots,y^{(32)}$ passed the verifier. We provide all 32 candidate proofs in a prompt, and ask the model to learn from the failed attempts and generate a new proof. We sample $k$ such generations, which constitutes one round of refinement. We do this for $T$ rounds, feeding in only the previous round’s generations. We use $k=16$ and $T=2$. In a small scale experiment we confirmed that this outperformed sampling $64$ candidates without refinement.
Medley Termination + Sorting Overall
Model (Easy) (Medium + Hard) (All)
GPT-4o (pass@32) 75.6% (65/86) 4.34% (5/115) 34.8% (70/201)
   + refinement 77.9% (67/86) 6.96% (8/115) 37.3% (75/201)
ntp-ctx-1.3B [37] 72.1% (62/86)   8.69% (10/115) 35.8% (72/201)
Table 3: miniCodeProps results. Number of specifications proven using full proof generation with GPT4 and best-first tactic generation with ntp-ctx-1.3B to the problem of verifying program specifications. The Medley section contains mostly of specifications that can be proven in several lines. Proofs of the sorting algorithm properties and termination lemmas are expected to require at least tens of lines and hours of programmer effort.
5 Results and Discussion